OpenStudy (anonymous):

WHICH FRACTION IS EQUIVALENT TO 5/6+7/8? A35/48 B6/7 C20/21 D41/24

OpenStudy (tkhunny):

Have you considered adding them? Common Denominator?

OpenStudy (jonnyvonny):

So, we should fine common denominators and solve this ourselves before looking at the answers. One common denominator is 48 (6*8), so lets do that. \[5/6+7/8=(5/6)(8/8)+(7/8)(6/6)=(40/48)+(42/48)=(82/48)\] Since this isn't re, we must reduce.\[1(34/48)=1(17/24)\] The answer is not there, which is odd, b/c even if you don't find a common denominator, you know that the fraction will exceed 1/1; it will have a while number.

OpenStudy (jonnyvonny):

Are you sure its addition and not subtraction?

OpenStudy (anonymous):

YEAH